40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jun 6 01:44:53 2009, rob , Update, Locking, ? 
    Reply  Sun Jun 7 04:57:39 2009, Yoichi, Update, Locking, ? 
       Reply  Tue Jun 9 23:25:24 2009, rob, Update, Locking, ? 
Message ID: 1663     Entry time: Tue Jun 9 23:25:24 2009     In reply to: 1660
Author: rob 
Type: Update 
Category: Locking 



Lock acquisition is proceeding smoothly for the most part, but there is a very consistent failure point near the end of the cm_step script.

Near the end of the procedure, while in RF common mode, the sensing for the MCL path of the common mode servo is transitioned from a REFL 166I signal which comes into the LSC whitening board from the demodulator, to another copy of the signal which has passed through the common mode board, and is coming out of the Length output of the common mode board.  We do this because the signal which comes through the CM board sees the switchable low-frequency boost filter, and so both paths of the CM servo (AO and MCL) can get that filter switched on at the same time.

The problem is occurring after this transition, which works reliably.  However, when the script tries to remove the final CARM offset, and bring the offset to zero, lock is abruptly lost.  DARM, CM, and the crossover all look stable, and no excess noise appears while looking at the DARM, CARM, MCF spectra.  But lock is always lost right about the same offset. 

Saturation somewhere?

 I've seen this before. At that time, the problem was gone spontaneously the next day.

You could stop just before the offset reaches zero and then try to slowly reduce the offset manually to see where is the threshold.



Well, it hasn't gone away yet.  It happened Sat, Mon, and Tues afternoon, as well as Friday.  The threshold varies slightly, but is always around ~200-300 cnts.   I've tried reducing the offset with the signal coming from the CM board and the signal not going through the CM board, I've also tried jumping the signal to zero (rather than a gradual reduction). 

Tonight we'll measure the MC length and set the modulation frequencies, and maybe try some MZ tweaking to do RFAMMon minimization.    

ELOG V3.1.3-